# Two Causally Related Needles in a Video Haystack

### v1

@@ 3.4 Quality Evaluation of Generated Questions
We utilize two state-of-the-art models not involved in the question generation process, ChatGPT-4.1 and Gemini-2.0-flash, to evaluate the quality of questions in Causal2Needles. The evaluation experiments are conducted from three perspectives: (1) Bridge entity accuracy, or if the bridge entity used for 2-needle question generation correctly connects the corresponding cause and effect events. In addition to the ground-truth bridge entity, we also construct a random baseline by randomly matching the bridge entities and the cause-effect event pairs. This baseline helps to verify that the evaluation LLMs do not always output high accuracy. (2) Semantic coherence, or whether the question is consistent with the story and does not introduce hallucination or content that contradicts the story. We use a 5-point scale, where 1 is the lowest score and 5 the highest. In the random baseline, we shuffle the correspondence between the questions and the cause-effect event pairs. (3) Readability of the questions, as reflected by the naturalness, grammar, and clarity, on a 5-point scale. Due to the difficulty in writing unreadable questions, we do not construct the random baseline for readability. The prompts for question evaluation are in Appendix Sec. F.
@@
Tab. 2 shows the evaluation results. The generated questions receive near-perfect scores on bridge entity accuracy and semantic coherence, and score highly on readability. Reassuringly, both random baselines score near the lowest possible, which is 0% for bridge entity accuracy and 1 for semantic coherence. These results indicate that the automatically generated questions have high quality and the evaluation processes is valid.
@@
Table 2: Evaluation results of generated questions. ‘1-N’, ‘VG 2-N’ and ‘ID 2-N’ denote 1-needle questions, visual grounding 2-needle questions and image description 2-needle questions, respectively. Numbers in parentheses indicate the performance of random baselines.
Models Bridge Entity (Random Baseline) Semantic Coherence Readability
1-N VG 2-N ID 2-N 1-N VG 2-N ID 2-N
ChatGPT-4.1 95.6% (0.3%) 4.62 (1.12) 4.99 (1.10) 4.74 (1.13) 4.85 4.83 4.67
Gemini-2.0-flash 95.0% (3.8%) 4.66 (1.02) 4.96 (1.01) 4.83 (1.00) 4.18 4.69 4.25
@@ 1-N VG 2-N ID 2-N 1-N VG 2-N ID 2-N
Table 3: Quantitative results (accuracy, %) of VLMs on our benchmark. ‘Forward’ refers to inputting video scenes in chronological order, while ‘Reverse’ uses reverse order. ‘Avg’ denotes results averaged over both orders. Best scores are in bold. ‘ID’ denotes image description and ‘VG’ denotes visual grounding.
Models 1-Needle Questions VG 2-Needle Questions ID 2-Needle Questions
@@
ChatGPT-4o 39.2 16.7 39.2 9.4 45.4 21.2 13.4 31.1 30.2 11.4 59.2
Gemini-1.5-pro 35.6 21.0 40.0 10.2 35.7 21.4 8.4 28.4 30.7 9.3 60.9
ChatGPT-4o-mini 33.4 17.4 22.9 5.0 32.4 11.9 5.2 24.9 17.4 5.1 52.3
Claude-3.5-sonnet 26.5 16.6 22.4 4.8 19.3 13.9 2.9 17.9 18.1 3.9 60.5
@@
LLaVA-OneVision-7B 18.0 4.6 14.7 0.0 17.0 5.6 0.1 10.8 10.2 0.1 28.3
LLaVA-Next-Video-7B 17.2 0.0 4.4 0.0 15.9 0.0 0.0 8.0 2.2 0.0 27.3
LongVA-7B 14.7 2.8 5.0 0.0 10.3 0.7 0.0 6.6 2.8 0.0 49.7
Qwen2.5VL-7B 13.6 27.6 17.7 5.0 11.2 18.9 1.9 19.4 18.3 3.4 43.2
LLaVA-Next-Video-34B 12.3 0.8 17.4 0.0 11.8 0.9 0.0 6.3 9.2 0.0 48.6
LongVU-7B 12.2 3.2 1.3 0.3 4.4 2.1 0.5 3.8 1.7 0.4 34.2
Aria-28B 12.1 19.0 14.8 0.6 18.7 18.1 0.1 18.9 16.5 0.4 43.0
InternVL2-8B 7.4 14.5 8.3 1.2 9.5 9.1 0.5 12.0 8.7 0.9 40.2
Random Chance 9.8 9.8 9.8 1.0 9.8 9.8 1.0 9.8 9.8 1.0 25.0
@@ 4.1 Evaluation Setup
We evaluate a total of 12 advanced VLMs, consisting of 4 proprietary models and 8 open-source models. Proprietary models include ChatGPT-4o (gpt-4o-2024-08-06) (OpenAI, 2024a), ChatGPT-4o-mini (gpt-4o-mini-2024-07-18) (OpenAI, 2024b), Gemini-1.5-pro (gemini-1.5-pro-002, Sep 2024) (Team et al., 2024), and Claude-3.5-sonnet (claude-3-5-sonnet-20241022) (Anthropic, 2024). Open-source models include LLaVA-Next-Video-7B (Zhang et al., 2024b), LLaVA-Next-Video-34B (Zhang et al., 2024b), LLaVA-OneVision-7B (Li et al., 2024a), LongVU-7B (Shen et al., 2024), Aria-28B (Li et al., 2024b), Qwen2.5VL-7B (Bai et al., 2025), LongVA-7B (Zhang et al., 2024a) and InternVL2-8B (Chen et al., 2024). Details about the input content and the chat template can be found in Appendix Sec. B.
Most models we tested do not natively support multiple video clips as input. As a workaround, we uniformly sample five frames from each clip and stacked them vertically as a single image. This design allows inputting multiple video clips without exceeding the image input limit of models, which is usually 32 frames. In our experiment, this trick yields the best performance. See Appendix Sec. C for details.
To prevent the model from using sentence indices as shortcuts, we input the full narration along with only a subset of the corresponding video clips. The narration and video clips are provided as separate, unaligned inputs—there is no sentence-to-clip pairing, which prevents the model from relying on structural alignment. The selected video clips span from the cause scene to the effect scene, with a random number of scenes added before and after this span. This setup further prevents shortcuts based on fixed answer location (i.e., the cause scene always appearing first and the effect scene last). Unless otherwise specified, a total of five extra scenes are padded across both ends.
@@ 4.2 Main Results
We present the performance of various VLMs on the 1-needle questions, Visual Grounding (VG) 2-needle questions, and Image Description (ID) 2-needle questions in Tab. 3. We report the accuracy on each type of questions, i.e., the proportion of correctly answered questions. For VG 2-needle questions, we separately compute the accuracy for Part 1, Part 2, and both parts answered correctly. In addition, we evaluate each model using both the original (forward) and reversed video input order, and report the average as the final result on VG 2-needle questions. The motivation of this design can be found in Sec. 4.3. We demonstrate the details of random chance calculation in Appendix Sec. B.
Figure 3: (a) The performance of models on VG 2-needle questions (Avg, Part 2) and 1-needle questions. The size of the dots indicates the model size. The stars indicate proprietary models. (b) The 2-needle visual grounding performance (Avg) on questions with different needle distances.
@@
Most models perform worse on Part 2 of VG 2-needle questions (Avg) than on 1-needle questions, despite both requiring cause clip identification. Fig. 3 (a) visualizes this pattern. Each point shows a model’s 1-needle performance (x-axis) and VG 2-needle Part 2 performance (y-axis). The diagonal line represents consistent performance across both question types, but most points fall below it, indicating that models strong on 1-needle questions often struggle with VG 2-needle questions. This highlights the deficiency of using only 1-needle questions to evaluate long-video understanding.
@@
In Fig. 3 (b), we present the VG 2-needle question performance (Avg) for ChatGPT-4o and Gemini-1.5-pro. Performance is measured on questions with varying distances between the cause and effect events. The results show a clear decline as distance increases, especially when the evaluation aggregates answers from both parts. This phenomenon indicates that the distance between multiple needles has a significant impact when joint understanding of these needles is required. More details are provided in Appendix Sec.C.
@@ 4.3 Pathological Behaviors of VLMs
As shown in Tab. 3, the accuracy of models responses varies depending on the positions of cause and effect scenes. In VG 2-needle questions, Part 1 is grounded in the effect video clip, while Part 2 is grounded in the cause clip. In the forward-ordered clip sequence, the cause clip appears earlier, and all proprietary models show higher accuracy in Part 2 compared to Part 1; for instance, the accuracy of ChatGPT-4o exhibits an accuracy increase of 22.5%. This result appears paradoxical, since from an information-processing perspective answering the Part 2 question requires an understanding of the video clip required for the Part 1 question. Why are models more accurate when locating the cause clip required by Part 2 than Part 1?
To answer this question, we reverse the order of the input video clips (the Reverse column of Tab. 3). In this setting, Part 1 achieves much higher accuracy than Part 2. This suggests that models pay more attention to scenes appearing earlier; therefore, visual grounding for the cause clips is easier under normal playback, while reverse playback facilitates the visual grounding of the effect clips. To mitigate the influence of position bias, we evaluate model accuracy of VG 2-needle questions using the average of forward and reverse clip ordering.
Figure 4: The answer distribution of various models in the forward evaluation of visual grounding 2-needle questions. “GT” denotes ground truth. “None” means no scene number is output.
@@
In Tab. 3, we observe that some open-source VLMs achieve extremely low scores, sometimes even zero, on VG 2-needle questions. When examined closely, we observe that some models tend to output the same prediction regardless of the question content. We call this phenomenon the static output bias.
In Fig. 4, we compare the output distributions of well-performing and poorly-performing models on VG 2-needle questions. For example, the ground truth of Part 1 is mainly in scene 5 and beyond, as the effect scene appears later. ChatGPT-4o shows a balanced distribution, aligning with its strong performance in Tab. 3. In contrast, the three open-source models predominantly select scene 1 as the answer, demonstrating severe output bias. Instead of considering the question’s content, they produce fixed responses. To mitigate this bias, we experiment with different prompting strategies, but the models consistently produce fixed responses, indicating an inherent flaw in these models. The test prompts and output patterns are detailed in Appendix Sec. F.
Table 4: Results of models on different input settings. Numbers in parentheses indicate the absolute increase over random chance. V: Video, N: Narration, Q: Question.
Models 1-Needle Questions Image Description Questions
@@ 4.4 Analysis of Dataset Bias
Textual bias refers to the extent that the dataset allows models to derive the correct answer from the accompanying text alone. To verify if our visual grounding test format effectively avoids textual bias, we test various models on 1-needle questions without providing the video clips. The results are presented in the first two columns of Tab. 5. After removing video input, model performance drops significantly to around random chance (9.8%) or lower. For instance, the performance of ChatGPT-4o drops from 39.2% to 3.8%. These results indicate that our setting effectively mitigates textual bias, ensuring a reliable evaluation of multimodal understanding abilities.

### v3

@@ 3.4 Quality Evaluation of Generated Questions
We conduct automatic evaluation and human evaluation of the quality of questions and bridge entities, which provide vital information for 2-needle questions. We evaluate 4 quality factors. Factor 1 is if the bridge entity is truly shared by the cause and effect events it is supposed to connect. We report the proportion of affirmative answers as the final score. Factor 2 is the correctness of purposely vague references to the bridge entities, or if the vague reference is indeed more ambiguous than the original bridge entity but still preserves its core meaning. The evaluation result for each vague reference should be “Yes” or “No”. We report the proportion of “Yes”. Factor 3 is the factual correctness of questions, or the extent to which a question is consistent with the story and does not introduce hallucination or contradiction to the story. We use a 5-point scale, where 1 is the lowest score and 5 the highest. Factor 4 is the readability of questions, as reflected by the naturalness, grammar, and clarity, on a 5-point scale.
To verify that the evaluation LLMs are not biased to always indicate high quality, we also create several random baselines. For Factor 1, we randomly match the bridge entities and the cause-effect event pairs, which should cause the LLM to answer No to shared existence of the bridge entity. For Factor 2, we randomly shuffle the correspondence between the bridge entities and the vague references. For Factor 3, we randomly shuffle the correspondence between the questions and the cause-effect event pairs. For Factor 4, due to the difficulty in writing unreadable questions, we do not construct any random baseline for readability.
We utilize two state-of-the-art models as the evaluation LLMs, ChatGPT-4.1 and Gemini-2.0-flash (neither is involved in the question generation process), and recruit five human annotators. The models evaluate both 1-needle and 2-needle questions, while the annotators evaluate 136 visual grounding 2-needle questions. For human evaluation, we adopt majority voting among the five annotators for Yes/No evaluations, and use the average for numerical evaluations.
@@
Tab. 2 shows the evaluation results. The generated questions receive near-perfect scores on all metrics. Reassuringly, all random baselines score near the lowest possible, which is 0% for shared existence of bridge entities and correctness of vague references, and 1 for factual correctness of questions. These results indicate that the automatically generated questions have high quality and the evaluation processes are valid. The prompts for question evaluation are in Appendix Sec. F.
@@
Table 2: Evaluation results of generated questions. VG and ID refer to visual grounding and image description, respectively. 1-N and 2-N denote 1-needle and 2-needle questions. Numbers in parentheses indicate the performance of random baselines.
Models
Shared Existence
of Bridge Entities
Correctness of
Vague References
Factual Correctness of Questions Readability of Questions
Noncausal
1-N
Causal
1-N
VG 2-N ID 2-N
Noncausal
1-N
Causal
1-N
VG 2-N ID 2-N
ChatGPT-4.1 95.6% (0.3%) 91.0% (3.6%) 4.71 (1.10) 4.62 (1.12) 4.99 (1.10) 4.74 (1.13) 4.91 4.85 4.83 4.67
Gemini-2.0-flash 95.0% (3.8%) 98.7% (2.5%) 4.75 (1.05) 4.66 (1.02) 4.96 (1.01) 4.83 (1.00) 4.75 4.18 4.69 4.25
Human 82.4% 98.5% - - 4.50 - - - 4.80 -
@@
Table 3: Quantitative results (accuracy, %) of VLMs on our benchmark. “Forward” refers to inputting video clips in chronological order, while “Reverse” uses reverse order. “Avg” denotes results averaged over both orders. Best scores are in bold.
Models
Noncausal 1-N
Questions
Causal 1-N
Questions
VG 2-N Questions
ID 2-N
Questions
@@
ChatGPT-4o 56.8 39.2 16.7 39.2 9.4 45.4 21.2 13.4 31.1 30.2 11.4 59.2
Gemini-1.5-pro 55.4 35.6 21.0 40.0 10.2 35.7 21.4 8.4 28.4 30.7 9.3 60.9
ChatGPT-4o-mini 39.9 33.4 17.4 22.9 5.0 32.4 11.9 5.2 24.9 17.4 5.1 52.3
Claude-3.5-sonnet 37.6 26.5 16.6 22.4 4.8 19.3 13.9 2.9 17.9 18.1 3.9 60.5
@@
Qwen2.5VL-32B 30.7 11.7 26.3 17.7 5.4 10.3 20.4 1.9 18.3 19.0 3.6 53.5
Qwen2.5VL-7B 17.5 13.6 27.6 17.7 5.0 11.2 18.9 1.9 19.4 18.3 3.4 43.2
LLaVA-Next-Video-34B 12.4 12.3 0.8 17.4 0.0 11.8 0.9 0.0 6.3 9.2 0.0 48.6
LLaVA-OneVision-7B 12.3 18.0 4.6 14.7 0.0 17.0 5.6 0.1 10.8 10.2 0.1 28.3
InternVL2-8B 11.6 7.4 14.5 8.3 1.2 9.5 9.1 0.5 12.0 8.7 0.9 40.2
LLaVA-Next-Video-7B 11.7 17.2 0.0 4.4 0.0 15.9 0.0 0.0 8.0 2.2 0.0 27.3
LongVA-7B 9.2 14.7 2.8 5.0 0.0 10.3 0.7 0.0 6.6 2.8 0.0 49.7
Aria-28B 7.0 12.1 19.0 14.8 0.6 18.7 18.1 0.1 18.9 16.5 0.4 43.0
LongVU-7B 3.3 12.2 3.2 1.3 0.3 4.4 2.1 0.5 3.8 1.7 0.4 34.2
Random Chance 9.8 9.8 9.8 9.8 1.0 9.8 9.8 1.0 9.8 9.8 1.0 25.0
@@ 4.1 Evaluation Setup
We evaluate a total of 13 advanced VLMs, consisting of 4 proprietary models and 9 open-source models. Proprietary models include ChatGPT-4o (gpt-4o-2024-08-06) (openai2024chatgpt4o), ChatGPT-4o-mini (gpt-4o-mini-2024-07-18) (openai2024chatgpt4o_mini), Gemini-1.5-pro (gemini-1.5-pro-002, Sep 2024) (team2024gemini), and Claude-3.5-sonnet (claude-3-5-sonnet-20241022) (anthropic2024claude3). Open-source models include LLaVA-Next-Video-7B (zhang2024video_llavanext), LLaVA-Next-Video-34B (zhang2024video_llavanext), LLaVA-OneVision-7B (li2024llava_onevision), LongVU-7B (shen2024longvu), Aria-28B (li2024aria), Qwen2.5VL-7B/32B (bai2025qwen2), LongVA-7B (zhang2024longva) and InternVL2-8B (chen2024far_internvl). To establish a human baseline, we employ two annotators to answer 316 randomly selected questions (including causal 1-needle, 2-needle visual grounding and image description). Details about the input content, the chat template of VLMs can be found in Appendix Sec. C.
Most models we tested do not natively support inputs containing multiple video clips. As a workaround, we uniformly sample five frames from each clip and stack them vertically as one input image. This allows us to stay within the image input limit, which is 32 images for most models. Compared to other possible techniques we experimented with, this trick yields the best performance (see Appendix Sec. D).
To prevent the models from using sentence indices as shortcuts, we input the full narration along with only a subset of the corresponding video clips. The narration and video clips are provided separately. We avoid aligning video clips with sentences in the input to prevent the models from relying on text to locate video clips. We input all video clips from the cause clip to the effect clip, as well as a total of five clip before and after this span; the numbers of clip before and after are random. This setup prevents the cause clip from always being the first clip and the effect clip always the last, eliminating shortcut learning based on location.
@@ 4.2 Main Results
We present the performance of various VLMs on the noncausal 1-needle questions, causal 1-needle questions, Visual Grounding (VG) 2-needle questions, and Image Description (ID) 2-needle questions in Tab. 3. We report the accuracy on each type of questions. For VG 2-needle questions, we separately compute the accuracy for Part 1, Part 2, and both parts answered correctly. In addition, we evaluate each model using both the original (forward) and reversed video clip order, and report the average of the two orders as the final result on VG 2-needle questions. This design is motivated by positional bias of the models, discussed in Sec. 4.3.
Figure 3: Comparison of performance of models on Part 2 of VG 2-needle questions (average order) and causal 1-needle questions. This is a fair comparison since both ask to retrieve the cause event. The dashed diagonal line represents equal performance across the two question types. Most points fall below the line, indicating that models are poorer at VG 2-needle questions than causal 1-needle questions. The size of the dots indicates the model size. The stars indicate proprietary models.
Figure 4: Performance on VG 2-needle questions as the distance between the two needle grows. We report the average-order performance for models and the forward-order performance for human annotators. The model performance declines whereas human performance stays mostly unchanged.
Causal Questions Are More Challenging Than Noncausal Questions.
From Tab. 3, we observe that most models perform substantially better on noncausal 1-needle questions than on causal ones. For instance, ChatGPT-4o and Qwen2.5-VL-32B achieve results that are 17.6% and 19.0% higher, respectively. This gap suggests that causal reasoning still poses a significant hurdle for current state-of-the-art models on long-video understanding.
@@
We compare the performance on Part 2 of VG 2-needle questions (average order) and causal 1-needle questions, since both ask for the retrieval of the cause event and constitute a fair comparison. Most models perform worse on the 2-needle questions than the 1-needle questions. Fig. 3 visualizes this pattern. This highlights the deficiency of using only 1-needle questions to evaluate long-video understanding.
@@
In Fig. 4 (a) and (b), we present the VG 2-needle question performance (average order) for ChatGPT-4o and Gemini-1.5-pro. Performance is measured on questions with varying distances between the cause and effect events. The results show a clear decline as distance increases, especially when the evaluation aggregates answers from both parts. This phenomenon indicates that the distance between multiple needles has a significant impact when joint understanding of these needles is required. More details are provided in Appendix Sec.D.
A possible confounder is the strength of causal relations [^1] We thank an anonymous reviewer for suggesting this.. That is, long distances between needles may be correlated with weak causal relationships. It may be the weak causality, rather than long distance, that causes performance decrease. To verify this conjecture, we test if humans can identify the cause clip over long distances. Two annotators answered 113 visual grounding 2-needle questions and the results with different needle distances are detailed in Fig. 4 (c). We can observe that for 24 samples with a needle distance between 6 and 8 clips, the human performance is 87.5% (Part 1), 87.5% (Part 2) and 87.5% (Both). This is even better than the performance on 97 samples with a distance between 3 and 5 clips: 83.5% (Part 1), 86.6% (Part 2) and 77.3% (Both). These results demonstrate that humans can easily undestand the causal relations. Thus, we argue that the model performance degradation mainly stems from the long distances between needles.
@@ 4.3 Pathological Behaviors of VLMs
As shown in Tab. 3, the accuracy of models responses varies depending on the positions of cause and effect clips. In VG 2-needle questions, Part 1 is grounded in the effect video clip, while Part 2 is grounded in the cause clip. In the forward-ordered clip sequence, the cause clip appears earlier, and all proprietary models show higher accuracy in Part 2 compared to Part 1; for instance, the accuracy of ChatGPT-4o exhibits an accuracy increase of 22.5%. This result appears paradoxical, since from an information-processing perspective answering the Part 2 question requires an understanding of the video clip required for the Part 1 question. Why are models more accurate when locating the cause clip required by Part 2 than Part 1?
To answer this question, we reverse the order of the input video clips (the Reverse column of Tab. 3). In this setting, Part 1 achieves much higher accuracy than Part 2. This suggests that models pay more attention to clips appearing earlier; therefore, visual grounding for the cause clips is easier under normal playback, while reverse playback facilitates the visual grounding of the effect clips. To mitigate the influence of position bias, we evaluate model accuracy of VG 2-needle questions using the average of forward and reverse clip ordering.
Figure 5: The answer distribution of various models in the forward evaluation of visual grounding 2-needle questions. GT denotes ground truth. None means no clip number is output. Predictions of opensource models are heavily concentrated in a few numbers, exhibiting significant bias.
@@
In Tab. 3, we observe that some open-source VLMs achieve extremely low scores, sometimes zero, on VG 2-needle questions. When examined closely, we observe that some models tend to output the same prediction regardless of the question content. We call this phenomenon the static output bias.
In Fig. 5, we compare the output distributions of well-performing and poorly-performing models on VG 2-needle questions. For example, the ground truth of Part 1 is mainly in clip 5 and beyond, as the effect clip appears after the cause. ChatGPT-4o shows a balanced distribution, aligning with its strong performance in Tab. 3. In contrast, the three open-source models predominantly select clip 1 as the answer, demonstrating severe output bias. We experiment with different prompting strategies as attempts to fix this bias, but the models consistently produce fixed responses, indicating an inherent flaw in these models. The test prompts and output patterns are in Appendix Sec. F.
Table 4: Results of models on different input settings. Numbers in parentheses indicate the absolute increase over random chance. 1-N: 1-needle, 2-N: 2-needle, V: Video, N: Narration, Q: Question. When V is removed in causal 1-N, performance decreases to near chance level, demonstrating minimal textual bias. When V+N are removed in ID 2-N, performances stay above chance level but are still far from saturation.
Models Causal 1-N Questions ID 2-N Questions
@@ 4.4 Analysis of Dataset Bias
Textual bias refers to the extent that the dataset allows models to derive the correct answer from the accompanying text alone. To verify if our visual grounding test format effectively avoids textual bias, we test various models on causal 1-needle questions without providing the video clips. The results are presented in the first two columns of Tab. 4. After removing video input, model performance drops significantly to around random chance (9.8%) or lower. For instance, the performance of ChatGPT-4o drops from 39.2% to 3.8%. These results indicate that our setting effectively mitigates textual bias, ensuring a reliable evaluation of multimodal understanding abilities.